# Lorentzian Graph Isomorphic Network

### v1

@@ Abstract
We introduce the Lorentzian Graph Isomorphic Network (LGIN), a novel graph neural network (GNN) designed to operate in hyperbolic spaces, leveraging the Lorentzian model to enhance graph representation learning. Existing GNNs primarily operate in Euclidean spaces, which can limit their ability to capture hierarchical and multi-relational structures inherent to complex graphs. LGIN addresses this by incorporating curvature-aware aggregation functions that preserve the Lorentzian metric tensor, ensuring embeddings remain constrained within the hyperbolic space by proposing a new update rule that effectively captures both local neighborhood interactions and global structural properties, enabling LGIN to distinguish non-isomorphic graphs with expressiveness at least as powerful as the Weisfeiler-Lehman test. Through extensive evaluation across nine benchmark datasets, including molecular and protein structures, LGIN consistently outperforms or matches state-of-the-art GNNs, demonstrating its robustness and efficacy in modeling complex graph structures. To the best of our knowledge, this is the first study to extend the concept of a powerful graph neural network to Riemannian manifolds, paving the way for future advancements in hyperbolic graph learning. The code for our paper can be found at https://github.com/Deceptrax123/LGIN.
@@ 1 Introduction
However, extending such frameworks to the hyperbolic analogy is inherently challenging due to the need to ensure that embeddings remain constrained within the Riemannian manifold. In particular, ensuring that the operations performed during aggregation and transformation preserve the hyperbolic metric is non-trivial. Motivated by the potential to unify high expressive power with geometry suited for hierarchical data, we propose the Lorentzian Graph Isomorphic Network (LGIN). LGIN operates on the hyperboloid (Lorentzian) model of hyperbolic space. It introduces a curvature-aware aggregation mechanism designed to maintain injectivity while effectively leveraging the negative curvature to capture both local neighborhood patterns and global graph structures. We show that LGIN enhances discriminative capability compared to Euclidean GINs, particularly for graphs with underlying hierarchical or complex relational patterns, by integrating structural isomorphism testing principles with the geometric inductive biases of hyperbolic space.
@@
  • We show that GNNs built using the hyperbolic-tangent-hyperbolic approach while preserving the metric tensor at the origin has the ability to distingush non-isomorphic graphs.
  • We propose Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets, as a more powerful alternative to Graph Isomorphic Network(GIN).
@@ 3.2 Definitions of Key Topological Concepts
Definition 1 (Differential Manifold).
@@
Definition 2 (Tangent Space).
@@
Definition 3 (Metric Tensor).
@@
Definition 6 (Parallel Transport).
@@ 4.1 Overview
Theorem 1.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two graphs. Consider a smooth Lorentzian manifold $(\mathcal{M},g)$ of constant negative curvature $c<0$, where $g$ is the Lorentzian metric tensor. Suppose we define an embedding
@@
such that each vertex $v\in V$ is mapped into the hyperboloid model of $\mathbb{H}^{n}$ If $G_{1}$ and $G_{2}$ are non-isomorphic, then their embeddings $\Phi(G_{1})$ and $\Phi(G_{2})$ cannot be related by a global isometry $f$ of $(\mathcal{M},g)$, i.e., there does not exist a diffeomorphism
@@
We present the proof for theorem 1 in the appendix section. Therefore, we can demonstrate that a graph neural network mapping node features onto a hyperboloid can achieve representational power comparable to the Weisfeiler-Lehman (WL) test. Further, the following must be taken into consideration:
  • A maximally powerful graph neural network requires both the neighborhood aggregation and readout functions to be injective(Xu et al., 2018).
@@ 4.2 Injectivity of Lorentz Transformation
To ensure the injectivity of the Lorentz transformation, we follow the hyperbolic-tangential-hyperbolic approach (Zhang et al., 2021a). This is because the tangent space at a point is locally isometric to Euclidean space. The Lorentz transformation used in our proposed method is given by:
@@
where $\mathbf{x}^{\mathcal{L}}\in\mathcal{L}_{c}^{n}$, $\mathbf{W}\in\mathbb{R}^{d\times n}$. This method ensures the first coordinate is consistently zero, signifying that the resultant transformation is invariably within the tangent space at $\mathbf{o}$. From (Amir et al., 2023), we have the following result for the Euclidean space: Consider shallow neural networks $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ of the form
@@
Therefore, a shallow(2-Layer) Lorentzian network based on transformation equation 14 with a pointwise Tanh/Sigmoid function is moment injective.
@@ 4.3 Lorentzian Graph Isomorphic Network
where $\mathbf{x}^{\mathcal{L}}$ is a point in the Lorentzian space, LT represents the Lorentz transformation which is a 2-layer model with Tanh activation, $k$ is the step and $T$ is the output of neighbor aggregation given by the cardinality aware Lorentz centroid method as follows:
@@
The pseudo code for the above rules is given in algorithm 1. Each variable is described in the appendix section.
@@ 5.2.2 Performance on Enzymes, Proteins and DD against state of the art graph neural networks
Table 2: Detailed Analysis on Proteins, Enzymes and DD. The best accuracy on the test splits are marked in bold while the second best score is underlined. Hyphen indicates the scores were not reported by the authors owing to out-of-memory(OOM) issues.
@@
Table 2 presents the detailed performance analysis across the Proteins, Enzymes, and DD datasets. We evaluate multiple models, including MLP, Vanilla GCN, GraphSage, MoNet, GAT, GatedGCN, GIN, RingGNN, 3WLGNN, and our proposed methods LGIN Variable and LGIN Fixed. On the Enzymes dataset, our LGIN Fixed model achieves the highest test accuracy of 71.50%, outperforming other models. The LGIN Variable variant follows closely with a test accuracy of 67.38%, showcasing the robustness of our method. For the DD dataset, GAT achieves the best test accuracy of 75.90%, surpassing other approaches. Our LGIN Variable model achieves a test accuracy of 73.73%, demonstrating its competitiveness and stability. In the Proteins dataset, our LGIN Variable model attains the highest accuracy at 79.10%, with the LGIN Fixed model following closely with 76.90%. This indicates the superior generalization capabilities of our proposed approach compared to existing models. Across all datasets, the proposed LGIN models show remarkable performance, consistently achieving either the best or second-best results, highlighting their ability to model complex graph structures effectively.
@@ 5.2.4 Empirical Analysis on Curvature and Parallel Transport
Figure 1: Abalation Study on Curvature and Parallel Transport on BBBP
@@
Appendix A Proof of Theorem 1
In this section, we present the proof to theorem 1. Before presenting the proof, we recall the theorem as
Theorem.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two graphs. Consider a smooth Lorentzian manifold $(\mathcal{M},g)$ of constant negative curvature $c<0$, where $g$ is the Lorentzian metric tensor. Suppose we define an embedding
$\Phi:\mathcal{G}\to\mathcal{M}$ (22)
such that each vertex $v\in V$ is mapped into the hyperboloid model of $\mathbb{H}^{n}$ If $G_{1}$ and $G_{2}$ are non-isomorphic, then their embeddings $\Phi(G_{1})$ and $\Phi(G_{2})$ cannot be related by a global isometry $f$ of $(\mathcal{M},g)$, i.e., there does not exist a diffeomorphism
$f:\mathcal{M}\to\mathcal{M}$ (23)
such that
$f^{*}g=g\quad\text{and}\quad f\circ\Phi(G_{1})=\Phi(G_{2})$ (24)
@@ Appendix A Proof of Theorem 1
Appendix B Variable Description of Algorithm 1

### v2

@@ Abstract
We introduce the Lorentzian Graph Isomorphic Network (LGIN), a novel graph neural network (GNN) designed to operate in hyperbolic spaces, leveraging the Lorentzian model to enhance graph representation learning. Existing GNNs primarily operate in Euclidean spaces, which can limit their ability to capture hierarchical and multi-relational structures inherent to complex graphs. LGIN addresses this by incorporating curvature-aware aggregation functions that preserve the Lorentzian metric tensor, ensuring embeddings remain constrained within the hyperbolic space by proposing a new update rule that effectively captures both local neighborhood interactions and global structural properties, enabling LGIN to distinguish non-isomorphic graphs with expressiveness atleast approximately as powerful as the Weisfeiler-Lehman test. Through extensive evaluations across nine benchmark datasets, including molecular and protein structures, LGIN consistently outperforms or matches state-of-the-art GNNs, demonstrating its robustness and efficacy in modeling complex graph structures. To the best of our knowledge, this is the first study to extend the concept of a powerful graph neural network to Riemannian manifolds, paving the way for future advancements in hyperbolic graph learning.
@@ 1 Introduction
However, extending such frameworks to the hyperbolic analogy is inherently challenging due to the need to ensure that embeddings remain constrained within the Riemannian manifold. In particular, ensuring that the operations performed during aggregation and transformation preserve the hyperbolic metric is non-trivial. Motivated by the potential to unify high expressive power with geometry suited for hierarchical data, we propose the Lorentzian Graph Isomorphic Network (LGIN). LGIN operates on the hyperboloid (Lorentzian) model of hyperbolic space. Our design prioritizes maintaining the intrinsic geometric properties and preserving the Lorentzian metric during aggregation and transformation, which presents challenges for achieving strictly injective message-passing. To address this, LGIN introduces a curvature-aware aggregation mechanism and update rule designed to approximate the expressive power of Euclidean GINs by carefully handling multisets within the hyperbolic geometry. We show that LGIN enhances discriminative capability compared to Euclidean GINs, particularly for graphs with underlying hierarchical or complex relational patterns, by integrating structural isomorphism testing principles with the geometric inductive biases of hyperbolic space by defining a theoretical framework for a powerfull GNN on hyperboli spaces.
@@
  • We show that GNNs built using the hyperbolic-tangent-hyperbolic approach while preserving the metric tensor at the origin high discriminative capability.
  • We propose the theoretical framework for Lorentzian Graph Isomorphic Netork(LGIN), with a modified cardinality-scaled aggregation rule to preserve the cardinality information of multisets, as an approximately powerful alternative to Graph Isomorphic Network(GIN) and other state of the art hyperbolic graoph neural networks such as LGCN.
@@ 3.2 Definitions of Key Topological Concepts
Definition 3.1 (Differential Manifold).
@@
Definition 3.2 (Tangent Space).
@@
Definition 3.3 (Metric Tensor).
@@
Definition 3.6 (Parallel Transport).
@@ 4.1 Overview
Theorem 4.1.
Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two graphs. Consider a smooth Lorentzian manifold $(\mathcal{M},g)$ of constant negative curvature $c<0$, where $g$ is the Lorentzian metric tensor. Suppose there exists an embedding
@@
that maps vertices of any graph $G\in\mathcal{G}$ into the hyperboloid model $\mathbb{H}^{n}$ and preserves graph structural properties (e.g., graph distance) through the Lorentzian distance on $\mathcal{M}$. If $G_{1}$ and $G_{2}$ are non-isomorphic, then their embeddings $\Phi(G_{1})$ and $\Phi(G_{2})$ cannot be related by a global isometry $f$ of $(\mathcal{M},g)$, i.e., there does not exist a diffeomorphism
@@
We present the proof for theorem 4.1 in appendix A. Therefore, we can demonstrate that a graph neural network mapping node features onto a hyperboloid can achieve representational power approximately comparable to the Weisfeiler-Lehman (WL) test. Further, the following must be taken into consideration:
  • A powerful graph neural networks requires a strictly injective aggregation and readout functions(Xu et al., 2018). However, directly translating strictly injective operations to Riemannian manifolds while preserving geometric properties is non-trivial. Our goal is to design a network that achieves high expressive power by carefully approximating these properties within the Lorentzian geometry.
@@ 4.2 Injectivity of Lorentz Transformation
We show the injectivity of the Lorentz Transformation on the manifold and through moment injectivity proposed by (Amir et al., 2023). To ensure the injectivity of the Lorentz transformation, we follow the hyperbolic-tangential-hyperbolic approach (Zhang et al., 2021a). This is because the tangent space at a point is locally isometric to Euclidean space. The Lorentz transformation used in our proposed method is given by:
@@
where $\mathbf{x}^{\mathcal{L}}\in\mathcal{L}_{c}^{n}$, $\mathbf{W}\in\mathbb{R}^{d\times n}$. This method ensures the first coordinate is consistently zero, signifying that the resultant transformation is invariably within the tangent space at $\mathbf{o}$. In the following sections. we prove the injectivity of the Lorentz transformation.
4.2.1 Injectivity on the Manifold
Proposition 4.2.
The Lorentzian transformation $\mathbf{W}\otimes_{c}^{\mathcal{L}}:\mathcal{H}^{n}\to\mathcal{H}^{d}$ defined by
$\mathbf{W}\otimes_{c}^{\mathcal{L}}x^{\mathcal{L}}=\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}x^{\mathcal{L}}\right)_{[1:n]}\right)$
is injective if and only if the weight matrix $\mathbf{W}\in\mathbb{R}^{d\times n}$ has full column rank, i.e., $\text{rank}(\mathbf{W})=n$. This requires $d\geq n$.
The proof of the above proposition is detailed in appendix B. Based on proposition 4.2, we ensure that the transformation matrix $\mathbf{W}$ has a full column rank by setting the number of output dimensions to be greater than the number of input dimensions for each layer.
4.2.2 Moment Injectivity
From (Amir et al., 2023), we have the following result for the Euclidean space: Consider shallow neural networks $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ of the form
@@ 4.2.2 Moment Injectivity
Therefore,we use a shallow(2-Layer) Lorentzian network based on transformation equation 14 with a pointwise Tanh/Sigmoid function and ensuring $\mathbf{W}$ has a full column rank.
@@ 4.3 Lorentzian Graph Isomorphic Network
where $\mathbf{x}^{\mathcal{L}}$ is a point in the Lorentzian space, LT represents the Lorentz transformation which is a 2-layer model with Tanh activation and output features greater than input, $k$ is the step and $T$ is the output of neighbor aggregation given by the cardinality aware Lorentz centroid method as follows:
@@
The pseudo code for the above rules is given in algorithm 1. Each variable and function is described in appendix C.
@@ 5.2.2 Performance on Enzymes, Proteins and DD against state of the art graph neural networks
Table 2: Detailed Analysis on Proteins, Enzymes and DD. The best accuracy on the test splits are marked in bold while the second best score is underlined. Hyphen indicates the scores were not reported by the authors. Hyperbolic GNN baselines have been marked with an asterisk(*).
@@
Table 2 presents the detailed performance analysis across the Proteins, Enzymes, and DD datasets. We evaluate multiple models, including Euclidean GNNs and hyperbolic GNN baselines(HGNN and H2H-GCN), alongside our proposed methods LGIN Variable and LGIN Fixed. On the Enzymes dataset, our LGIN Fixed model achieves the highest test accuracy of 71.500%, significantly outperforming the hyperbolic baselines HGNN(51.3%) and H2H-GCN(61.3%), and other Euclidean models. GAT achieves the second best performance (68.500%). The LGIN Variable variant follows closely with 67.380%. For the DD dataset, the hyperbolic baseline H2H-GCN achieves the highest test accuracy of 78.2%. GAT is the second best at 75.900%. HGNN performs very close to GAT (75.8%). Our LGIN Variable model achieves a test accuracy of 73.730%, performing comparably to HGNN and better than many Euclidean baselines, and is the fourth best overall on this dataset. LGIN Fixed has lower performance on DD (70.300%). In the Proteins dataset, our LGIN Variable model attains the highest accuracy at 79.100%, surpassing both LGIN Fixed (76.900%) and the hyperbolic baselines HGNN(73.7%) and H2H-GCN(74.4%). H2H-GCN achieves the third best performance on this dataset. Overall, the results demonstrate that LGIN models remain highly competitive across these datasets. LGIN Fixed is the top performer on Enzymes, while LGIN Variable leads on Proteins. Although H2H-GCN is the best on DD, LGIN Variable and Fixed show strong performance on Proteins and Enzymes respectively, highlighting LGIN’s ability to model complex graph structures effectively within the Lorentzian geometry, competing well against both Euclidean and other hyperbolic methods.
@@ 5.2.3 Performance on MoleculeNet Datasets
Figure 1: Abalation Study on Curvature Type and Parallel Transport on BBBP
@@ 5.2.4 Empirical Analysis on Curvature and Parallel Transport
Appendix A Proof of Theorem 4.1
@@
Appendix B Proof of Proposition 4.2
Proof.
Let the transformation be denoted by $F(x^{\mathcal{L}})=\mathbf{W}\otimes_{c}^{\mathcal{L}}x^{\mathcal{L}}$. We can decompose $F$ into a sequence of maps: $F=f_{5}\circ f_{4}\circ f_{3}\circ f_{2}\circ f_{1}$, where:
  1. 1.
    $f_{1}:\mathcal{H}^{n}\to T_{o}\mathcal{H}^{n}$, $f_{1}(x^{\mathcal{L}})=\log_{o}^{c}(x^{\mathcal{L}})$. The logarithmic map centered at the origin $o$ of the hyperboloid model $\mathcal{H}^{n}$ is a global diffeomorphism from $\mathcal{H}^{n}$ (the upper sheet) to $T_{o}\mathcal{H}^{n}$. As a diffeomorphism, $f_{1}$ is bijective and thus injective.
  2. 2.
    $f_{2}:T_{o}\mathcal{H}^{n}\to\mathbb{R}^{n}$, $f_{2}({\bm{c}}{v})={\bm{c}}{v}_{[1:n]}$. In the hyperboloid model embedded in $\mathbb{R}^{n+1}$, the tangent space at the origin $o=(1,0,\dots,0)$ consists of vectors ${\bm{c}}{v}=(v_{0},v_{1},\dots,v_{n})\in\mathbb{R}^{n+1}$ such that $\langle o,{\bm{c}}{v}\rangle_{\eta}=-1\cdot v_{0}+\sum_{i=1}^{n}0\cdot v_{i}=-v_{0}=0$. Thus $T_{o}\mathcal{H}^{n}=\{(0,v_{1},\dots,v_{n})\mid v_{i}\in\mathbb{R}\}$. The map $f_{2}$ extracts the spatial components $(v_{1},\dots,v_{n})$, providing an isomorphism between $T_{o}\mathcal{H}^{n}$ and $\mathbb{R}^{n}$. As an isomorphism, $f_{2}$ is bijective and thus injective.
  3. 3.
    $f_{3}:\mathbb{R}^{n}\to\mathbb{R}^{d}$, $f_{3}({\bm{c}}{y})=\mathbf{W}{\bm{c}}{y}$. This is a standard linear transformation. A linear transformation $f_{3}$ is injective if and only if its null space, $\ker(\mathbf{W})=\{{\bm{c}}{y}\in\mathbb{R}^{n}\mid\mathbf{W}{\bm{c}}{y}={\bm{c}}{0}\}$, contains only the zero vector. By the Rank-Nullity Theorem ($n=\dim(\ker(\mathbf{W}))+\text{rank}(\mathbf{W})$), this is equivalent to $\text{rank}(\mathbf{W})=n$. The maximum possible rank for $\mathbf{W}\in\mathbb{R}^{d\times n}$ is $\min(d,n)$. Thus, having rank $n$ requires $d\geq n$. $f_{3}$ is injective if and only if $\mathbf{W}$ has full column rank ($n$).
  4. 4.
    $f_{4}:\mathbb{R}^{d}\to T_{o}\mathcal{H}^{d}$, $f_{4}({\bm{c}}{z})=(0,{\bm{c}}{z})$. Similar to $f_{2}$, this map takes a vector ${\bm{c}}{z}\in\mathbb{R}^{d}$ and forms a vector $(0,z_{1},\dots,z_{d})\in\mathbb{R}^{d+1}$, which lies in $T_{o}\mathcal{H}^{d}$. This map is an isomorphism between $\mathbb{R}^{d}$ and $T_{o}\mathcal{H}^{d}$. As an isomorphism, $f_{4}$ is bijective and thus injective.
  5. 5.
    $f_{5}:T_{o}\mathcal{H}^{d}\to\mathcal{H}^{d}$, $f_{5}({\bm{c}}{u})=\exp_{o}^{c}({\bm{c}}{u})$. The exponential map centered at the origin $o$ is a global diffeomorphism from $T_{o}\mathcal{H}^{d}$ to $\mathcal{H}^{d}$ (the upper sheet). As a diffeomorphism, $f_{5}$ is bijective and thus injective.
The composite function $F=f_{5}\circ f_{4}\circ f_{3}\circ f_{2}\circ f_{1}$ is injective if and only if all functions in the composition are injective. We have shown that $f_{1},f_{2},f_{4},$ and $f_{5}$ are injective. Therefore, $F$ is injective if and only if $f_{3}$ is injective.
As shown in point 3, $f_{3}$ is injective if and only if the matrix $\mathbf{W}\in\mathbb{R}^{d\times n}$ has full column rank ($n$). This implies that the output dimension $d$ must be greater than or equal to the input dimension $n$ ($d\geq n$).
Thus, the Lorentzian transformation defined in eq. 14 is injective if and only if $\mathbf{W}$ has full column rank.
To demonstrate this explicitly, assume $F(x_{1}^{\mathcal{L}})=F(x_{2}^{\mathcal{L}})$ for $x_{1}^{\mathcal{L}},x_{2}^{\mathcal{L}}\in\mathcal{H}^{n}$.
$\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}\right)=\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}\right)$
Since $f_{5}=\exp_{o}^{c}$ is injective, we have:
$\left(0,\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}\right)=\left(0,\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}\right)$
Equating the spatial components (or applying $f_{4}^{-1}$ which is also injective):
$\mathbf{W}\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}=\mathbf{W}\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$
Let ${\bm{c}}{y}_{1}=\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}$ and ${\bm{c}}{y}_{2}=\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$. Both ${\bm{c}}{y}_{1},{\bm{c}}{y}_{2}\in\mathbb{R}^{n}$. The equation is $\mathbf{W}{\bm{c}}{y}_{1}=\mathbf{W}{\bm{c}}{y}_{2}$. If $\mathbf{W}$ has full column rank, the linear map $f_{3}({\bm{c}}{y})=\mathbf{W}{\bm{c}}{y}$ is injective, which implies ${\bm{c}}{y}_{1}={\bm{c}}{y}_{2}$.
$\left(\log_{o}^{c}x_{1}^{\mathcal{L}}\right)_{[1:n]}=\left(\log_{o}^{c}x_{2}^{\mathcal{L}}\right)_{[1:n]}$
Since $f_{2}^{-1}$ maps $\mathbb{R}^{n}$ back to $T_{o}\mathcal{H}^{n}$ injectively, this implies the equality of the tangent vectors:
$\log_{o}^{c}x_{1}^{\mathcal{L}}=\log_{o}^{c}x_{2}^{\mathcal{L}}$
Finally, since $f_{1}=\log_{o}^{c}$ is injective, this implies:
$x_{1}^{\mathcal{L}}=x_{2}^{\mathcal{L}}$
Thus, if $\mathbf{W}$ has full column rank, $F(x_{1}^{\mathcal{L}})=F(x_{2}^{\mathcal{L}})$ implies $x_{1}^{\mathcal{L}}=x_{2}^{\mathcal{L}}$, proving $F$ is injective.
Conversely, assume $\mathbf{W}$ does not have full column rank. Since $\mathbf{W}\in\mathbb{R}^{d\times n}$, this means $\text{rank}(\mathbf{W})<n$. By the Rank-Nullity Theorem, $\dim(\ker(\mathbf{W}))=n-\text{rank}(\mathbf{W})>0$. This means there exists a non-zero vector ${\bm{c}}{v}\in\mathbb{R}^{n}$ such that $\mathbf{W}{\bm{c}}{v}={\bm{c}}{0}$. Let ${\bm{c}}{u}=f_{2}^{-1}({\bm{c}}{v})\in T_{o}\mathcal{H}^{n}$. Since $f_{2}^{-1}$ is injective and ${\bm{c}}{v}\neq{\bm{c}}{0}$, ${\bm{c}}{u}\neq{\bm{c}}{0}$. Let $x^{\mathcal{L}}=f_{1}^{-1}({\bm{c}}{u})\in\mathcal{H}^{n}$. Since $f_{1}^{-1}$ is injective and ${\bm{c}}{u}\neq{\bm{c}}{0}$, $x^{\mathcal{L}}\neq o$. Consider $F(x^{\mathcal{L}})$ and $F(o)$. $F(x^{\mathcal{L}})=f_{5}(f_{4}(f_{3}(f_{2}(f_{1}(x^{\mathcal{L}})))))=f_{5}(f_{4}(f_{3}({\bm{c}}{u})))=f_{5}(f_{4}(\mathbf{W}{\bm{c}}{v}))=f_{5}(f_{4}({\bm{c}}{0}))$. $f_{4}({\bm{c}}{0})=(0,{\bm{c}}{0})\in T_{o}\mathcal{H}^{d}$. $F(x^{\mathcal{L}})=\exp_{o}^{c}(0,{\bm{c}}{0})$. The exponential map of the zero vector in the tangent space at $p$ is always $p$. So, $\exp_{o}^{c}(0,{\bm{c}}{0})=o\in\mathcal{H}^{d}$. Now consider $F(o)$: $F(o)=\exp_{o}^{c}\left(0,\mathbf{W}\left(\log_{o}^{c}o\right)_{[1:n]}\right)$. $\log_{o}^{c}o$ is the zero vector in $T_{o}\mathcal{H}^{n}$, which is $(0,{\bm{c}}{0})\in\mathbb{R}^{n+1}$. $\left(\log_{o}^{c}o\right)_{[1:n]}={\bm{c}}{0}\in\mathbb{R}^{n}$. $F(o)=\exp_{o}^{c}(0,\mathbf{W}{\bm{c}}{0})=\exp_{o}^{c}(0,{\bm{c}}{0})=o\in\mathcal{H}^{d}$. So $F(x^{\mathcal{L}})=o$ and $F(o)=o$. Since $x^{\mathcal{L}}\neq o$, we have found two distinct points $x^{\mathcal{L}},o\in\mathcal{H}^{n}$ that map to the same point $o\in\mathcal{H}^{d}$. Thus, $F$ is not injective if $\mathbf{W}$ does not have full column rank.
Therefore, the Lorentzian transformation $\mathbf{W}\otimes_{c}^{\mathcal{L}}$ is injective if and only if the matrix $\mathbf{W}$ has full column rank. ∎
Appendix C Variable Description of Algorithm 1